Given the following assumptions, develop a cost estimate (Labor only) that integrates all of the cost elements into a final bid price to the customer. What is the total cost for the project? How long will the project take to complete?
Function points in the project: 260
Software engineers assigned to this team: 4
Function point productivity per software engineer: 5 per month
Workdays per typical month: 22
Productive hours per typical workday: 8
Gross hourly wage rate per software engineer (does not include fringe benefits): $50 per hour
Overhead (fringe benefit, other direct overhead) rate 42%
G&A rate: 11.5%
Proposed Profit: 15%
Answer #1

Q. How long will the project take to complete?

Solution:-

Function points in the project = 260

Points per month per software engineer = 5

number of software engineers = 4

So we can say that the total points in a month will be 5*4=20

Time taken in completion of the project (In months) = total function points of the project / total point per month

= 260/20=13 month

Q. What is the total cost for the project?

Particulars Quantity / Amount ($)
Hours / day 8
Workdays / Month 22
Wage rate / software engineer $ 50
Total Hours in a month 176
Total wage / month / engineer $ 8800
Total engineers 4
Total monthly wages $ 35200
total months work will be done 13
total wages in 13 month $ 457600
overhead rate of the total wages 42%
overhead cost $ 192192
G&A rate of the total wages 11.50%
G&A Cost $ 52624
Total cost of the porject $ 702416

Solution

The final bid price to the customer can be given including the profit in the total cost

Total cost of the project = $702416

Proposed profit = 15% = 15/100*702416 = $105362

Total Proposed bid to the customer = Total cost + proposed profit

= $ 702416 + $ 105362 = $ 807778